Calculate number of neutrons present in 12x10^25 atoms Of oxygen (80^17) :
(Given: NA = 6x 10^23)
(a) 1800
(b) 1600
(C) 1800 NA
(d) 3200 NA
1
Expert's answer
2020-11-11T08:57:54-0500

Solution: The answer is "1800N_A" . i.e option C.


Atomic number "=P=8" .


Atomic Mass "= N+P = 17."


Therefore "N = 9," Which is number of neutrons.


Total number of neutrons in "12\times10^{25} atoms = 9\times 12\times 10^{25}=\frac{9\times12\times10^{25}\times N_A}{6\times10^{23}}"


"=1800N_A" .